# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v2

@@ IV-A Spinal cord and brainstem module validation
Figure 5: (a) Average tracking error using single mode and two modes across three trajectories. (b) Average weight generated by the thalamus module across three trajectories.
IV-B Thalamus and cerebellum module validation
Fig. 5a shows that single gravity compensation modes from the cerebellum module are only effective within narrow load ranges (e.g., the light mode around 1 kg and the heavy mode around 3.5 kg), with errors otherwise increasing nearly linearly with load variation. In contrast, the thalamus module adaptively synthesizes both modes, maintaining stable joint error across the entire 1–3.5 kg range, thereby validating its design. As shown in Fig. 5b, reinforcement learning enables the thalamus module to adjust mode weights according to load mass, assigning a smaller weight to the light mode and a greater weight to the heavy mode as the load increases. This adaptive mechanism mirrors human motor control, where movements are achieved through linear combinations of motor primitives[sanger2000human], thus highlighting the biomimetic nature of the thalamus and cerebellum module.
@@ IV-B Thalamus and cerebellum module validation
Figure 6: Average joint error across the three trajectories under varying load conditions during operation. (a) 0-1.5s. (b) 12-14s. (c) 24-25.5s. (d) 0-1.5s without the brainstem module. (e) 0-1.5s without the thalamus module. (f) 0-1.5s without the cerebellum module.
IV-C Framework validation
Fig. 6 illustrates the average trajectory tracking error across three trajectories under different load conditions. As shown in Fig. 6a, when a 1.5kg load is applied at 0s, the algorithm converges to a stable state within approximately 1.5s; after 12s of operation, adding an additional 2kg load again leads to convergence within about 2s; and upon removing the added load at 24s, convergence is also achieved within 2s. In contrast, the cerebellum-inspired network proposed in [abadia2019robot, abadia2021cerebellar] requires nearly 100s to adapt to a 500g change in load, highlighting the superior dynamic performance and convergence speed of our framework. Furthermore, we observe that, compared with cold-start initialization, subsequent load changes during ongoing operation induce smaller error fluctuations, as reflected by reduced peak magnitude at 12s and 24s, which can be attributed to the membrane potential dynamics of spiking neurons that preserve historical information and facilitate smoother adaptation.
Fig. 6b, 6c and 6d further show that the convergence process consistently exhibits three distinct error peaks following each load change. These peaks directly correspond to the hierarchical loop structure of the framework: the spinal cord module (first-order level) updates at the control frequency and therefore corresponds to every point on the curve; the brainstem and thalamus modules (second-order level) yield the earlier two peaks which reflect the learning convergence process; and the cerebellum module (third-order level) produces the latest peak owing to its longest processing cycle. The temporal separation of these peaks validates the layered organization of the control framework.
@@ IV-C Framework validation
To verify this correspondence, ablation experiments were conducted by selectively disabling individual modules, as shown in Fig. 6e, 6f and 6g. Removing the brainstem’s learning capability by fixing feedback control coefficients reduces the number of peaks to two and delays the first peak, indicating diminished adaptability to load change. Similarly, disabling the thalamus module by fixing mode-combination weights eliminates its characteristic peak, leaving only brainstem and cerebellar responses. Finally, replacing the cerebellum’s adaptive torque generation with fixed compensation suppresses the cerebellar peak and weakens the thalamic one, since the thalamus relies on cerebellar outputs to optimize weight assignment. These results confirm the functional contributions of each module and their hierarchical cooperation.
In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence. This step-by-step activation phenomenon mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
V Experiment results
This section will present real-world experiments on a robotic arm platform to demonstrate contribution (4). All spiking neurons, except those in the spinal cord module, are implemented in Python using an open-source SNN learning framework, while the spinal cord module is realized in C++ through direct modeling of membrane potential dynamics. The deep learning framework, SNN framework, and computing hardware are identical to that used in the simulation. The robotic arm platform is a Flexiv Rizon 4s equipped with a Flexiv GRAV gripper, as shown in Fig. 7.
@@ V Experiment results
The experiments are conducted under nine load conditions, including two material types (sand-filled and water-filled containers) and five mass levels ranging from 1.5 kg (gripper with empty container) to 3.5 kg (container with 2 kg load). The upper limit is chosen below the 4 kg load capacity of the robotic platform to maintain a safety margin.
Eight end-effector trajectories (T1–T8, defined in the Appendix) are tested, all with vertically downward orientation and origin at the base flange center. Among them, T1–T5 represent slower, lower-dynamic trajectories, while T6–T8 correspond to faster, higher-dynamic ones. For statistical validity, each trajectory–load–controller combination is repeated 10 times. In all experiments presented in this section, the trajectory tracking effect is evaluated using the RMSE of the end-effector position and orientation. At each time step, the position and orientation RMSE are defined as
@@
where $x_{d},y_{d},z_{d},\theta_{d},\psi_{d},\phi_{d}$ denote the desired end-effector position in Cartesian space and the desired pitch, yaw, and roll angles, respectively, and $x,y,z,\theta,\psi,\phi$ denote the corresponding actual values. The average position and orientation RMSE of a trajectory are defined as
$\text{RMSE}_{p}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{p}(t)$ (31)
$\text{RMSE}_{o}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{o}(t)$ (32)
where $N$ represents the trajectory length.
@@
Figure 8: Trajectory tracking result under different trajectory and load conditions. (a) Position RMSE under sand load. (b) Orientation RMSE under sand load. (c) Position RMSE under water load. (d) Orientation RMSE under water load.
@@
Fig. 8a and 8b show the trajectory tracking errors of the robotic arm when the end-effector container is filled with sand, under various trajectory and load conditions. Red and blue bars represent the proposed method and the factory-installed industrial-grade Flexiv Cartesian Position Control, respectively. Cross-hatched and slanted-hatched bars indicate the error ranges of trajectories T1–T5 (lower dynamic group) and T6–T8 (higher dynamic group) for a given load, while error bars denote the extrema of the errors. Four observations can be drawn:
@@
1. Higher tracking accuracy: For any load mass, the proposed method yields smaller mean position and orientation errors, as indicated by the consistently lower red bars compared with the blue bars. Quantitatively, the Flexiv method achieves average errors of $1.789\times 10^{-2}$ m (position) and $1.685\times 10^{-2}$ rad (orientation) across all trajectories and load conditions, whereas the proposed method achieves $1.447\times 10^{-2}$ m and $1.473\times 10^{-2}$ rad, corresponding to reductions of 19.1% and 12.6%, respectively. These results demonstrate that the proposed method achieves higher control accuracy, thereby exhibiting superiority in the “Precision” indicator of agility.
2. Adaptation to load variation: As load mass increases, position errors of the proposed method gradually increase while orientation errors decrease, and both remain generally lower than those of the Flexiv method, as shown by the upward trend of red bars in Fig. 8a and the downward trend in Fig. 8b. This observation is consistent with the simulation results, confirming that the proposed framework can adapt to arbitrary load masses within a practical range, validating the design of the cerebellar and thalamic modules.
@@
3. Consistency across trajectories: Within each trajectory group of similar dynamics, the proposed method exhibits greater consistency across different trajectory shapes, as the red bars for both hatching styles are shorter than their blue counterparts. Specifically, the variation ranges of Flexiv errors in position and orientation are $0.447\times 10^{-2}$ m and $0.592\times 10^{-2}$ rad, while those of the proposed method are $0.152\times 10^{-2}$ m and $0.363\times 10^{-2}$ rad, corresponding to reductions of 66.0% and 38.7%. This aligns with simulation findings, indicating that the proposed method can rapidly adapt online to arbitrary motion patterns, validating the brainstem and spinal cord modules and demonstrating an advantage in the agility indicator of “Transitions between diverse motion patterns”.
4. Performance under higher dynamics: For the higher-dynamic group, the performance improvement of the proposed method is more pronounced, as the gap between red and blue slanted bars is larger than that between the cross-hatched bars. Quantitatively, in the higher-dynamic group, the proposed method reduces position and orientation errors by 24.7% and 17.0% relative to the Flexiv method, exceeding the average improvements reported in conclusion 1. This outcome is expected, since the Flexiv method, being model-based, necessarily involves linear simplifications to ensure computational efficiency. While such simplifications suffice at low speeds where nonlinear effects (e.g., Coriolis and centrifugal forces) are weak, they deteriorate at higher speeds. In contrast, the proposed method adapts online to nonlinear dynamics, leading to smaller error increases. These results suggest that the proposed framework is more suitable for fast motions, demonstrating superiority in the agility indicator of “Speed” .
Fig. 8c and 8d present results when the container is filled with water. Compared to sand, water introduces stronger dynamic variations, thereby exerting greater influence on control performance. The proposed method yields mean errors of $1.442\times 10^{-2}$ m (position) and $1.519\times 10^{-2}$ rad (orientation), representing increases of 0.9% and 2.0% relative to the sand condition; the Flexiv method yields $1.779\times 10^{-2}$ m and $1.681\times 10^{-2}$ rad, representing decreases of 1.2% and 0.7%, respectively. Both methods show comparable trends across different load materials, indicating the ability to adapt to varying material properties. Moreover, the proposed method consistently demonstrates similar control performance under sand and water conditions, and analogous conclusions to the four points above can be drawn, which are omitted here for brevity.
In summary, the experiment results demonstrate that the proposed framework consistently outperforms the industrial-grade control algorithm across the three key indicators of agility—precision, adaptability to diverse motion patterns, and speed as mentioned in section I—thereby highlighting its effectiveness and the enhanced agility capacity provided by the brain-inspired design.
VI Conclusion and future works
In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments. The framework introduces a hierarchical architecture with five modules implemented using SNNs, and has been validated through both simulations and real-world robotic experiments. Results show that this framework operates in a manner similar to the human CNS and can outperform an industrial-grade baseline. Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
[Experiment trajectory setting]
@@ VI Conclusion and future works
Yanbo Pang received the B.Eng. from Sun Yat-sen University in 2022 and the M.Sc. from Tsinghua University in 2025. He is currently an intern at Booster Robotics. His research interests include neuromorphic computing, neurorobotics and robotic arm control.
Qingkai Li received his B.S. from Tsinghua University in 2021. He is currently a Ph.D. candidate in the Department of Automation at Tsinghua University, with research interests spanning neuromorphic computing, bio-inspired algorithms, imitation learning, and whole-body control.
Mingguo Zhao received the B.Eng., M.Sc. and Ph.D. from Harbin Institute of Technology in 1995, 1997 and 2001, respectively. He is currently a Full Professor with the Department of Automation and a director of the Beijing Innovation Center for Future Chips, Tsinghua University. His research interests include brain-inspired robot control and general robotic platform based on neuromorphic computing and hardware.

### v3

@@ IV-A Module design validation
Figure 4: Performance of the cerebellum and thalamus modules. (a) Average RMSE across three trajectories using single motor pattern and two motor patterns. (b) Average motor pattern weight coefficients across three trajectories generated by the thalamus module.
Figure 5: Performance of the brainstem module.
Fig. 5 illustrates the evolution of the average stiffness coefficients $k_{p,i},i=1,\cdots,7$ and damping coefficients $k_{v,i},i=1,\cdots,7$ for each joint across three trajectories. The results show that the brainstem module can adaptively generate feedback control gains for the spinal cord module. Notably, the stiffness and damping coefficients of joints 4 and 6 generally exhibit opposite trends. Quantitative analysis confirms this observation: the correlation coefficients between $k_{p,4}$ and $k_{v,4}$, and between $k_{p,6}$ and $k_{v,6}$, are 0.5640-0.5640 and 0.8515-0.8515, respectively. This behavior is consistent with human motor control. Physiological studies have shown that joint stiffness and damping in the human arm are inversely correlated [5], indicating that the brainstem module exhibits biomimetic characteristics. Overall, these results demonstrate that the design of the brainstem module satisfies the proposed requirements.
IV-B Resilience performance evaluation
@@
Figure 6: Resilience performance under unstructured load changes. (a) Evolution of average RMSE across the three trajectories. (b) Expanded view of 0-1.5 s. (c) Expanded view of 12-14 s. (d) Expanded view of 24-25.5 s. (e) Expanded view of 0-1.5 s without the brainstem module. (f) Expanded view of 0-1.5 s without the thalamus module. (g) Expanded view of 0-1.5 s without the cerebellum module.
Fig. 6 illustrates the resilience of the proposed framework under unstructured load variations. As shown in Fig. 6a, when a 1.5 kg load is applied at $t=0$ s, the framework recovers after a learning period of approximately 1.5 s. After 12 s of operation, an additional 2 kg load is introduced, and the system again recovers within about 2 s. When the added load is removed at 24 s, recovery is achieved within approximately 1.5 s. In contrast, cerebellum-inspired networks [2, 3, 1] require nearly 100 s to adapt to a 500 g load change, while CBMC-V2 [32] requires about 15 s to converge after initialization.
Furthermore, the resilience performance improves over time. Although the load change at 12 s is larger than that at 0 s, the resulting tracking error fluctuation during recovery is considerably smaller, as shown in Fig. 6b and Fig. 6c. Similarly, as illustrated in Fig. 6d, the fluctuation induced by load removal at 24 s is even smaller, despite the mass change being identical to that at 12 s. In contrast, the error fluctuations caused by load variations in [2, 3, 1] do not diminish as operation time increases.
It can also be observed from Fig. 6b, Fig. 6c, and Fig. 6d that the recovery process following each load change consistently exhibits three distinct error peaks. These peaks indicate that the framework converges in a hierarchical manner. Specifically, the spinal cord module (first-level) operates at the real-time control frequency and corresponds to the response reflected at each data point of the curve; the brainstem and thalamus modules (second-level) account for the first two peaks; and the cerebellum module (third-level) is responsible for the final peak. This stepwise activation pattern is analogous to human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [15].
@@
To verify that these peaks are associated with the framework’s hierarchical structure, ablation experiments are conducted by selectively disabling individual modules. As shown in Fig. 6e, removing the learning capability of the brainstem module by fixing the feedback control coefficients reduces the number of peaks to two and delays the first peak. Similarly, as illustrated in Fig. 6f, disabling the thalamus module by fixing the motor pattern weighting coefficients eliminates the second peak, leaving only the responses associated with the brainstem and cerebellum modules. Finally, as shown in Fig. 6g, replacing the cerebellum network output with a fixed compensation torque suppresses the cerebellum module peak and weakens the thalamus module response, since it relies on cerebellum module outputs to optimize motor pattern integration.
In summary, these results demonstrate that the proposed framework exhibits strong and progressively improving resilience during operation. This capability arises from the synergistic interaction of its hierarchical structure and modular design. Removing any module, control level, or pathway leads to a noticeable degradation in performance.
V Practical results
This section will present practical experiment results on a commercial robotic arm platform to demonstrate the precision and adaptation performance. The robotic arm platform is a Flexiv Rizon 4s equipped with a Flexiv GRAV gripper, as shown in Fig. 7. The deep learning framework, SNN framework, control cycle of each level, and computing hardware are identical to that used in the simulation, and the Flexiv RDK version is 1.5.1.
@@ V Practical results
To represent different tasks, five load masses, two load materials, and eight end-effector trajectories are designed. The load masses range from 1.5 kg (the gripper with an empty container) to 3.5 kg (the container with a 2 kg load), with increments of 500 g. The two load materials are sand and water. All trajectories are designed with a vertically downward orientation, with their definition and parameters provided in the Appendix. Among them, T1–T5 represent slower trajectories, whereas T6–T8 correspond to faster trajectories. The framework is evaluated on each trajectory ten times under different combinations of load mass and material.
In all experiments presented in this section, the trajectory tracking error is evaluated using the RMSE of the end-effector position and orientation. At each time step, the position and orientation RMSE are defined as
@@
where $x_{d},y_{d},z_{d},\theta_{d},\psi_{d},\phi_{d}$ denote the desired end-effector position in Cartesian space and the desired pitch, yaw, and roll angles, respectively, and $x,y,z,\theta,\psi,\phi$ denote the actual position in Cartesian space and the actual pitch, yaw, and roll angles, respectively. The average position and orientation RMSE of a trajectory are defined as
$\text{RMSE}_{p}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}_{p}(t)$ (31)
$\text{RMSE}_{o}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}_{o}(t)$ (32)
where $L$ represents the trajectory length.
@@
Figure 8: Precision and adaptation performance under different trajectory and load settings. (a) Position RMSE under sand load. (b) Orientation RMSE under sand load. (c) Position RMSE under water load. (d) Orientation RMSE under water load.
@@
Fig. 8 shows the trajectory tracking errors of the robotic arm under various trajectory and load conditions. Fig. 8a-b correspond to sand loads, while Fig. 8c-d correspond to water loads. The red and blue bars represent the proposed method and the factory-installed Flexiv Cartesian Position Control, respectively. Each bar indicates the range of average errors obtained by repeating each trajectory ten times under a given load condition, while the error bars denote the overall range of errors across all trajectories in the corresponding group. The slanted-hatched bars correspond to the fast trajectory group (T1–T5), whereas the cross-hatched bars represent the slow trajectory group (T6–T8). Five observations can be drawn:
@@
1. Precision: For all load masses, the proposed method consistently yields smaller average position and orientation errors, as indicated by the uniformly lower red bars compared with the blue bars. Under sand-load conditions, the Flexiv controller achieves average errors of $1.800\times 10^{-2}$ m and $1.694\times 10^{-2}$ rad across all trajectory and mass configurations, whereas the proposed method reduces these errors to $1.430\times 10^{-2}$ m and $1.489\times 10^{-2}$ rad, corresponding to reductions of 20.6% and 12.1%, respectively. Under water-load conditions, the Flexiv controller yields average errors of $1.779\times 10^{-2}$ m and $1.681\times 10^{-2}$ rad, while the proposed method achieves $1.442\times 10^{-2}$ m and $1.519\times 10^{-2}$ rad, resulting in reductions of 18.9% and 9.6%, respectively.
@@
2. Adaptation to trajectory shape: Within each trajectory group, the proposed method exhibits greater consistency across different trajectory shapes, as the red bars under both hatching styles are consistently shorter than their blue counterparts. For sand-load conditions, the average variation ranges of the mean position and orientation errors for the Flexiv method are $0.441\times 10^{-2}$ m and $0.596\times 10^{-2}$ rad across all mass configurations, whereas those of the proposed method are reduced to $0.146\times 10^{-2}$ m and $0.393\times 10^{-2}$ rad, corresponding to reductions of 66.9% and 34.1%. For water-load conditions, the Flexiv method exhibits average variation ranges of $0.455\times 10^{-2}$ m and $0.588\times 10^{-2}$ rad, while the proposed method achieves $0.157\times 10^{-2}$ m and $0.338\times 10^{-2}$ rad, corresponding to reductions of 65.5% and 39.4%, respectively.
3. Adaptation to trajectory speed: The proposed method produces smaller tracking errors for both slow and fast trajectory groups. Notably, the improvement is more pronounced for the fast trajectories, as evidenced by the larger gap between the red and blue slanted-hatched bars compared with that between the cross-hatched bars. Under sand-load conditions, for the fast trajectory group, the proposed method reduces position and orientation errors by 29.4% and 17.5%, respectively, relative to the Flexiv method—exceeding the average improvements reported in Observation 1. Under water-load conditions, corresponding reductions of 28.5% and 14.1% are achieved.
4. Adaptation to load mass: As the load mass increases, the position error of the proposed method increases, while the orientation error decreases; nevertheless, both remain consistently lower than those of the Flexiv method. This behavior is reflected by the upward trends of the red bars in Fig. 8a and Fig. 8c, and the downward trends in Fig. 8b and Fig. 8d.
5. Adaptation to load material: Water introduces stronger dynamic variations due to its complex hydrodynamic behavior, thereby exerting a greater influence on control performance. Nevertheless, similar trends are observed when comparing Fig. 8c–d with Fig. 8a–b, and all of the above observations consistently apply to both load materials.
Taken together, the experimental results demonstrate that the proposed framework consistently outperforms the factory-installed method across a wide range of task conditions, including variations in trajectory shape, execution speed, load mass, and load material. The framework achieves higher precision and stronger adaptation to unstructured changes, with particularly advantages in high-speed scenarios.
VI Conclusion and future work
In this paper, we propose a CNS-inspired, SNN-based motion control framework aimed at achieving agile manipulation in unstructured environments. The framework adopts a hierarchical architecture comprising five modules, three control levels, and two pathways, and is validated through both simulation studies and practical experiments. Taken together, the results demonstrate that the proposed method operates in a manner analogous to the human CNS and fulfills three out of the four key requirements of agile motion control outlined in Section I, namely precision, adaptation, and resilience.
Energy efficiency is not explicitly analyzed in this paper, as it depends on multiple factors, including the deep learning and SNN frameworks, computational hardware, and the controlled robotic arm platform. In addition, orientation tracking performance is consistently inferior to position tracking performance, indicating room for further improvement. Moreover, adaptation to varying load masses can be further enhanced, as tracking precision degrades under changes in load mass. These aspects, along with a more thorough analysis of velocity tracking error, will be investigated in future work.
[Practical experiment trajectory setting]